OpenStudy (anonymous):

The coordinates of the centre of the circle with equation x^2-6x+y^2-2y=18 is? Working out please ! :D

OpenStudy (anonymous):

(3,1)

OpenStudy (anonymous):

(x-3)^2-9+(y-1)^2-1=18 (x-3)^2+(y-3)^2=28

OpenStudy (anonymous):

Is the equation Centre is (3,1) like electic engineer said

OpenStudy (anonymous):

Sorry (x-3)^2+(y-1)^2=28

OpenStudy (anonymous):

do you know how to complete a square?

OpenStudy (anonymous):

Yeah i just remembered lol

OpenStudy (anonymous):

6x+(3/2)^2 and so on..... i have bad memory :(